### v1

# The reliability of hybrid functionals for accurate fundamental and optical gap prediction of bulk solids and surfaces

Abstract
Hybrid functionals have been considered insufficiently reliable for the prediction of band gaps in solids and surfaces. We revisit this issue with a new generation of optimally-tuned range-separated hybrid functionals, focusing on the reconstructed Si(111)-(2$\times$1) and Ge(111)-(2$\times$1) surfaces. We show that certain hybrid functionals can accurately predict the surface-state and bulk fundamental and optical gaps, as well as projected band structures of these surfaces, by combining ground-state and time-dependent density functional theory.
Density functional theory (DFT) has become a standard method to determine the electronic structure of both molecular and extended systems Parr and Yang (1989); Dreizler and Gross (1990); Martin (2008); Burke (2012); Teale et al. (2022). In practice, it is often used within the Kohn-Sham (KS) framework Kohn and Sham (1965), where the original, many-electron system is mapped onto an equivalent system of non-interacting electrons. This KS mapping is exact in principle, in the sense of obtaining the same density as that of the original system. However, differences between the lowest unoccupied and highest occupied KS eigenvalues almost always underestimate the true fundamental band gap of the system, even if the exact functional is used Perdew and Levy (1983); Sham and Schlüter (1983). Most modern DFT calculations and codes use this KS scheme in practice with (semi-)local approximations to the exchange-correlation (XC) functional, which leads to severe failures in the prediction of optical properties of solids when used within time-dependent (TD) DFT Maitra (2016).
Due to the current limitations of standard TDDFT, ab initio many-body perturbation theory approaches Hedin (1965); Onida et al. (2002) have become a popular choice for the computation of electronic and optical excitations in solids. Specifically, calculation of charged excitations within the $GW$ approximation Hybertsen and Louie (1986), followed by calculations of neutral excitations using the Bethe-Salpeter equation (BSE) Rohlfing and Louie (1998); Albrecht et al. (1998) are considered state-of-the-art Martin et al. (2016). But the computationally costly ‘one-shot’ $G_{0}W_{0}$ and the subsequent $G_{0}W_{0}$-BSE (dubbed $GW$-BSE in this letter), which are often used in practice to obtain optical absorption spectra, rely on the underlying density functional used. This starting point dependence is a known deficiency and an ongoing topic of research Rinke et al. (2005); Fuchs et al. (2007); Chen and Pasquarello (2014); Leppert et al. (2019); Gant et al. (2022); Golze et al. (2019). One solution to the limitations of KS theory is to use the generalized KS (GKS) scheme, which maps the original system onto one of partially interacting particles that are still represented by a single Slater determinant, thereby allowing the introduction of non-multiplicative XC potential operators Seidl et al. (1996). Within this scheme, more advanced approximations, notably meta-generalized gradient approximations (MGGA) Cangi et al. (2019); Sun et al. (2015); Patra et al. (2019); Traoré et al. (2019); Yang et al. (2016); Lebeda et al. (2024, 2023); Aschebrock and Kümmel (2019) and hybrid functionals  Becke (1993); Perdew et al. (1996a, 2017); Shimazaki and Asai (2009); Skone et al. (2014, 2016); Sun et al. (2020); Jana et al. (2020); Miceli et al. (2018a); Yang et al. (2022), have emerged as serious contenders for the solution of the band gap problem in solids.
An area in which the accuracy of advanced GKS-based functionals remains an important open issue is that of a functional that can consistently and accurately predict not only the fundamental gap, but also the optoelectronic properties of bulk materials, and more complex cases, like surfaces Jain et al. (2011). In particular, Jain et al. Jain et al. (2011) examined the performance of popular hybrid functionals such as PBE0 Perdew et al. (1996a), HSE Heyd et al. (2003, 2006), and B3LYP Stephens et al. (1994) for the Si(111)-(2$\times$1) surfaces. They concluded that none can be used in a ground-state DFT calculation to consistently and accurately predict the fundamental gap, $E_{g}$, and the optical gap, $E_{opt}$ (quantities that differ by the exciton binding energy), of both bulks and surfaces. This task remains challenging even today, given that many advanced per-orbital correction methods that significantly improve the prediction of $E_{g}$ have yet to to be applied to optical excitations Mahler et al. (2022); Yang et al. (2022); Nguyen et al. (2018); Dabo et al. (2014); Ma and Wang (2016).
Here, we reexamine the possibility of obtaining both bulk and surface $E_{g}$ and $E_{opt}$ using the above-mentioned hybrid functionals, as well as a new class of Wannier optimally-tuned screened range-separated hybrid functionals (WOT-SRSH) Wing et al. (2021) that are applicable to 3D solids, including layered materials Ohad et al. (2022, 2023); Camarasa-Gómez et al. (2024); Sagredo et al. (2024). We show that it is now possible to meet this challenge and predict accurate and reliable gaps for bulks and surfaces using slab calculations with certain hybrid functionals. In particular using WOT-SRSH, and for the standard semiconductors studied here also HSE, using ground-state DFT and TDDFT for charged and neutral excitations, respectively.
The generalized Kohn-Sham (GKS) equation for a screened range-separated hybrid (SRSH) functional can be defined as Refaely-Abramson et al. (2013)
$\begin{split}\bigg{[}-\frac{\bm{\nabla}^{2}}{2}+V(\bm{r})+V_{H}([n];\bm{r})+V_{X}^{SRSH}([n];\bm{r})+\\ V_{SLc}([n],\bm{r})\Bigg{]}\phi_{i}(\bm{r})=\epsilon_{i}\phi_{i}(\bm{r})\end{split}$ (1)
with
$\begin{split}V_{X}^{SRSH}=\alpha\hat{V}_{F}^{SR,\gamma}+(1-\alpha)V_{SLx}^{SR}+\\ \frac{1}{\epsilon_{\infty}}\hat{V}_{F}^{LR,\gamma}+\bigg{(}1-\frac{1}{\epsilon_{\infty}}\bigg{)}V_{SLx}^{LR}.\end{split}$ (2)
Here $V,V_{H},V_{X}^{SRSH}$, and $V_{c}$ are the external, Hartree, exchange, and correlation potentials, respectively. $\hat{V}_{F}$ is the Fock potential operator, SL denotes semi-local exchange or correlation, SR and LR correspond to short or long range potentials, and $\epsilon_{i}$ and $\phi_{i}$ are the generalized KS eigen energies and associated orbitals. Finally, $\epsilon_{\infty}$ corresponds to the orientally averaged, clamped-ion dielectric constant, and $\gamma$ is the range separation parameter.
A key aspect of range-separated hybrid functionals, which differentiates them from global hybrids (e.g. PBE0 or B3LYP), is the different fraction of Fock exchange in the treatment of short- and long-range exchange associated with the $1/r$ Coulomb interaction, controlled by the $\gamma$ parameter Leininger et al. (1997). We note that global hybrids are special cases of Eq. (1), e.g., PBE0 Perdew et al. (1996a) is obtained when $\alpha=1/\epsilon_{\infty}=1/4$, independent of $\gamma$. Similarly, HSE can be recovered for $\alpha=1/4$ and $\gamma=0.2{\rm\AA}^{-1}$, with $\epsilon_{\infty}\rightarrow\infty$. Different choices of $\alpha,\gamma,\epsilon_{\infty}$ lead to different hybrid functionals, and must be made judiciously. For example, it has been shown that using an ‘arbitrary amount’ of Fock exchange $\alpha$ can lead to qualitatively incorrect defect physics in 2D materials Chen et al. (2022).
Table 1: Fundamental bulk and surface state gaps obtained from Si(111)-(2$\times$1) and Ge(111)-(2$\times$1), and compared to reference $GW$ and experiments.
(eV) Si(111)2$\times$1 Ge(111)2$\times$1
Bulk Surface Bulk Surface
PBE 0.61 0.37 0.07 0.39
AM05 0.50 0.40 0.08 0.43
SCAN 0.84 0.45 0.18 0.50
$\alpha$-r2SCAN 1.72 0.77 1.37 1.80
PBE0 1.78 1.00 1.47 1.20
HSE 1.16 0.64 0.83 0.74
SRSH(($\alpha,\gamma$)Wing et al. (2021)) 1.15 0.64 0.86 0.75
Reference
$GW$ 1.23 Rohlfing and Louie (1999) 0.69 Ciccacci et al. (1986) 0.72 Aulbur et al. (2000) 0.67 Zhu and Louie (1991)
Exp 1.17 Madelung (2004) 0.75 Ciccacci et al. (1986) 0.75 Aulbur et al. (2000) 0.65 Nicholls and Reihl (1989)
For tuned screened range-separated hybrid (SRSH) functionals, the choice of parameters is flexible, but not arbitrary. Typically, $\alpha$ is chosen along with $\gamma$ to satisfy the ionization potential (IP) theorem (i.e. the optimally-tuned [OT-]SRSH Refaely-Abramson et al. (2013)), or an IP ansatz associated with removal of localized charge (i.e. the Wannier-localized optimally-tuned [WOT-]SRSH Wing et al. (2021)). In other words, this ansatz serves as our physical constraint. By satisfying this known exact condition, using $\epsilon_{\infty}$ to constrain the long-range Coulomb interaction, and using an image-charge correction (when removing a charge from the Wannier function), the WOT-SRSH functional has shown improvements of the fundamental bands gaps, band structures and optical gaps of materials Wing et al. (2021); Ohad et al. (2023, 2022); Sagredo et al. (2024); Camarasa-Gómez et al. (2024).
The WOT-SRSH functional is part of a recently emerging class of functionals  Kronik et al. (2012); Tal et al. (2020); Stein et al. (2010); Refaely-Abramson et al. (2013); Autschbach and Srebro (2014); Miceli et al. (2018b). A distinct aspect of this method is the ansatz used in the localization scheme for the wavefunction of extended systems, introduced in Ref. Ma and Wang (2016), specifically using maximally-localized Wannier functions. We note that HSE is also a range-separated hybrid (RSH) functional, albeit a special one because it lacks long-range exchange, and thus in general lacks the correct asymptotic potential. Therefore, while there is merit to exploring modifications in the HSE parameters Moussa et al. (2012), it is unsuitable for the systematic tuning procedure presented above.
Even given the success of WOT-SRSH for bulk systems, surfaces pose additional challenges to the tuning procedure. A major part of the added difficulty comes from the fact that the slab structures, used to simulate surfaces, are environments with a different dielectric screening than in the bulk. Therefore limits that can typically be taken for the dielectric medium in 2D systems, (particularly $\epsilon_{\infty}\rightarrow 1$ Qiu et al. (2016); Cudazzo et al. (2011); Andersen et al. (2015)), including the generalized tuning procedures for 2D systems based on them Camarasa-Gómez et al. (2024), cannot be used in surfaces. A natural question then arises: Is it possible to use tuned ($\alpha,\gamma$) parameters obtained for the bulk, even for surface-containing slabs? In other words, how transferable are the tuned ($\alpha,\gamma$) parameters outside of the ideal bulk scenario? More generally, how reliable are various hybrid functionals in this context?
(a)Refer to caption (b)Refer to caption
(c,d)Refer to caption
Figure 1: Projected band structure for slabs containing the reconstructed (a) Si(111) and (b) Ge(111)-(2$\times$1) surfaces (see SI for details on structures used). We show the band structure of the whole system with blue solid line, and highlight the contribution of the surface states with pink (Si) and green (Ge) dots that are proportional to their weight in the wavefunction, obtained using the SRSH functional with tuned bulk parameters. (c) Linear absorption spectrum, Im$[(\epsilon_{XX}+\epsilon_{YY}+\epsilon_{ZZ})/3]$, for Si(111)-(2$\times$1) and (d) Ge(111)-(2$\times$1), of the lowest-energy bright excitations associated with each surface, with a 0.1 eV broadening. Reported experimental values are denoted by the vertical gray bar at 0.47 eV Macfarlane et al. (1957) (0.49 eV Rohlfing and Louie (2000)), and previous $GW$-BSE Jain et al. (2011); Rohlfing and Louie (2000) results are given by the blue vertical line at 0.43 eV(0.51 eV) for Si (and Ge). The purple, magenta, and green lines are the results of the TDDFT calculations with SRSH, HSE, and PBE0, respectively, for the reconstructed Si (and Ge) (111)- (2$\times$1) surface represented by the slab. See SI for detailed discussion of Lorentzian broadening value used to average the plotted peaks, as well as analysis of asymmetry of the lowest peak of the HSE Ge spectrum.
We answer the above questions using two well-known reconstructed semiconductor surfaces Srivastava (1997): Si(111)-(2$\times$1) Pandey (1981); Reining and Del Sole (1991); Northrup et al. (1991), and Ge(111)-(2$\times$1) Rohlfing and Louie (2000). We model these surfaces with slab structures that contain properties of both bulk and surfaces. Results for these systems can be readily compared to literature experimental Uhrberg et al. (1982); Perfetti et al. (1987); Ciccacci et al. (1986); Macfarlane et al. (1957); Nannarone et al. (1980); Aulbur et al. (2000) and theoretical Jain et al. (2011); Rohlfing and Louie (2000); Aulbur et al. (2000) values for bulk and surface fundamental and optical gaps.
The (2$\times$1) reconstruction is produced by taking an ideal Si (or Ge) (111) surface and pushing two of the surface atoms toward each other. After surface relaxation, $p_{z}$ orbitals form $\pi$ bonded chains along the [001] direction of the surface. The dangling $p_{z}$ orbitals in turn, generate two surface-state bands inside the bulk-forbidden gap Srivastava (1999), yielding distinctly different surface-state and bulk gaps of this material. For compatibility with past $GW$ results, we use the known positive buckling configuration for the Si case and the negative buckling for Ge  Nie et al. (2004).
All calculations are performed using the Vienna Ab initio Simulation Package VASP Kresse and Furthmüller (1996, 1996), on a 48 atom unit cell (24 layers), with 20 Å of vacuum. More computational details, including assessment of slab size, structural information, convergence tests, and comparison to results obtain with the QUANTUM ESPRESSO (QE) package Giannozzi et al. (2009), are found in the supplementary information (SI) Sup . We emphasize that for the optical spectra we are only concerned with the peak associated with the surface state, which is the lowest-energy one for both Si and Ge slabs. The peaks correspond to the lowest excitations with non-zero oscillator strengths. Our reported gaps are an average of the lowest bright excitons with a broadening of $0.1$ eV. See the SI Sup for more details and analysis. Bulk optical peaks are known to appear at higher energies and are well-prediced by TDDFT with SRSH for both Si and Ge Wing et al. (2019, 2020).
As a first step, we calculate the surface projected band structures, using the bulk-tuned Wing et al. (2021) WOT-SRSH functional. In Fig. 1 (a) and (b), we plot the projected surface states of the Si and Ge surfaces in pink and green, respectively. Clearly, the bands of the surface states can be seen within the bulk bands of each material, as described above. More projected band structures calculated with other functionals can be found in the SI Sup . We find good agreement with previously computed projected band structures Jain et al. (2011); Zhu and Louie (1991) for these surfaces. The fundamental gap for the surface states occurs at J. We observe a finite band gap at $\Gamma$, for Ge, whereas a negligible gap is found with standard semilocal or meta-GGAs tested for comparison (see SI).
In Table LABEL:table1, we present calculations for the bulk and surface state fundamental gaps for Si(111)-(2$\times$1) and Ge(111)-(2$\times$1) using PBE Perdew et al. (1996b), PBE0 Perdew et al. (1996a), HSE06 Krukau et al. (2006); Heyd et al. (2006), AM05 Armiento and Mattsson (2005), SCAN Sun et al. (2015), $\alpha$-r2SCAN (i.e. r2SCAN with Fock exchange) DelloStritto et al. (2023). We further compare with previous $GW$ and $GW$-BSE calculations Jain et al. (2011); Rohlfing and Louie (2000). We only investigate some currently popular meta-GGAs, but we make note of the recent success of newer functionals in similar bulk systems Aschebrock and Kümmel (2019); Lebeda et al. (2023, 2024). We find the expected underestimations of the bulk and surface state gaps of the Si structure using the semi-local PBE functional, by $0.56$ and $0.38$ eV, respectively, when compared to experiment (as does AM05 with an absolute error (AE) of 0.67 and 0.35 eV, respectively). The global hybrid PBE0 overestimates both the bulk and surface state gap by $0.61$ and $0.25$ eV, respectively. The HSE06 hybrid functional performs with good accuracy, with an AE of $0.11$ and $0.01$ eV for the surface state and bulk gap, respectively. The meta-GGA SCAN underestimates the bulk gap by $0.33$ eV, and the surface state gap by $0.30$ eV. The SRSH functional, using previously tuned Wing et al. (2021) Si bulk ($\alpha$, $\gamma$) pairs from the WOT-SRSH functional, performs on par with HSE06, with an AE of $0.02$ and $0.13$ eV, for bulk and surface state gap respectively. Finally, using the recommended $\alpha=0.15$ Fock exchange DelloStritto et al. (2023) the r2SCAN Furness et al. (2020) functional does get good agreement for the surface state fundamental gap, within 0.02 eV, but it overestimates the bulk gap by 0.55 eV with respect to the experimental value. Summarizing, the SRSH and HSE functionals outperform all other density approximations and are within 0.08 and 0.07 eV of previous $GW$ results Rohlfing and Louie (1999); Ciccacci et al. (1986); Aulbur et al. (2000); Zhu and Louie (1991), respectively.
We repeat this set of calculations with a similar slab representing the reconstructed Ge(111)-(2$\times$1) surface (see right side of table LABEL:table1). We find the same general trends as in the Si case, but with a few caveats: First, as shown above, the fundamental gap of bulk Ge at $\Gamma$ is known to be almost zero for most semilocal and meta-GGA functionals. Second, even when using a hybrid, specifically HSE, as a starting point for the $GW$ calculations, an overestimation of the bulk Ge gap is seen. This shows the sometimes inconsistent nature of ‘one-shot’ $GW$, even with a hybrid functional starting point used in previous studies Rinke et al. (2005); Gant et al. (2022); Golze et al. (2019).
Finally, using linear response TDDFT  Casida (1996); Tretiak and Chernyak (2003), within the Tamm-Dancoff approximation, it is possible to obtain the linear absorption spectra of the reconstructed surfaces, here in arbitrary units for the oscillator strength. The results of our calculations are seen in Fig. 1 for the reconstructed Si (c) and Ge (d) slabs for PBE0, HSE, and SRSH functionals. The attainment of the optical gap for these slabs, made clear by Fig. 1, goes beyond the expectations of Ref. Jain et al. (2011) from hybrid functionals, owing to the use of TDDFT rather than ground-state DFT. We find that both the SRSH and HSE functionals perform with good accuracy, and get within 0.08 and and 0.14 eV of the experimental optical gap for Si respectively (for Ge the errors are 0.13 eV and 0.16 eV, respectively). When compared to previously reported $GW$-BSE optical gaps for the Si case, the SRSH and HSE are within 0.12 and 0.1 eV (for Ge those errors are 0.15 and 0.14 eV) respectively. PBE0 tends to overestimate the optical gap for the Si and Ge surface. While the calculated spectra in Fig. 1 allow for a visual comparison of the TDDFT excited states associated with the surface states for different functionals, the magnitude of the peak depends on the supercell and slab size, and therefore are taken as arbitrary. The lowest-energy optical peaks for the Si(111)-(2$\times$1) and Ge(111)-(2$\times$1) slabs, at the WOT-SRSH level, are well within the $\pm$ 0.2 eV experimental uncertainty and prior GW-BSE calculations Jain et al. (2011). In the SISup , we report the individual excitons that make up the absorbance peaks and analyze the effects of Lorenzian broadening. Other studies have shown that in-gap states induce novel low-lying excitons not present in the bulk, e.g., those originating with a mixture of bulk states and surface states Hernangómez-Pérez et al. (2023). A detailed study of surface excitons would be an interesting subject for future work.
To conclude, we find that it is possible to reliably predict both the bulk and surface state fundamental gaps, of Si(111)-(2$\times$1) and Ge(111)-(2$\times$1) with the SRSH (using previously tuned bulk $\alpha$, and $\gamma$ parameters) and HSE functionals, with good accuracy. More importantly, we find that it is possible to obtain accurate optical gaps, using these hybrid functionals, with TDDFT calculations. While both SRSH and HSE functionals perform equally well here, we expect to see deviations in the performance of larger gap materials, similar to the bulk case (e.g. Fig. 3 in SI of Wing et al. (2021), and Chen et al. (2018)). Previous work more than a decade ago Jain et al. (2011) concluded that hybrid functionals (at the time) did not give reliable band gaps for bulk systems or surfaces, and that fortuitously obtaining a good surface state and bulk gap with a hybrid functional did not guarantee the optical gap with ground state DFT calculation. Here, we overcome this by using the DFT and TDDFT frameworks with a new generation of SRSH, namely WOT-SRSH, where $\alpha$ and $\gamma$ are constrained by an IP ansatz, can lead to bulk and surface fundamental and optical gaps in excellent agreement with experiments and previous $GW$-BSE results. Finally, while a tuning procedure is not yet available for surfaces, the reliable performance seen with the transferability of bulk parameters for these reconstructed surfaces, provides an indication that the construction of this class of hybrids goes in the correct direction.
I Acknowledgements
F.S. acknowledges discussions with Prof. M. Jain which clarified important aspects of past work. M.C.-G. and F.S. thank Prof. A. Ramasubramaniam for fruitful discussions. M.C.-G. is grateful to the Azrieli foundation for the award of an International Postdoctoral Fellowship. L.K. acknowledges additional support from the Aryeh and Mintzi Katzman Professorial Chair and from the Helen and Martin Kimmel Award for Innovative Investigation. The formalism used in this work was funded through NSF–Binational Science Foundation Grant No. DMR-2015991 and the Israel Science Foundation; and our calculations were supported by the Theory of Materials Program at LBNL, funded by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, under Contract No. DE-AC02-05CH11231. Computational resources were provided by the NSF funded ACCESS program (frontera supercomputer),at the Texas Advanced Computing Center (TACC), the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility at Lawrence Berkeley National Laboratory, and the Chemfarm facility at the Weizmann Institute of Science.

### v2

@@
 The reliability of hybrid functionals for accurate fundamental and optical gap prediction of bulk solids and surfaces
+Abstract
 Abstract
@@
 We find the same general trends as in the Si case, but with a few caveats: First, as shown above, the fundamental gap of bulk Ge at $\Gamma$ is known to be almost zero for most semilocal and meta-GGA functionals.
-Second, even when using a hybrid, specifically HSE, as a starting point for the $GW$ calculations, an overestimation of the bulk Ge gap is seen.
+Second, even when using a hybrid, specifically HSE, as a starting point for the $GW$ calculations, an overestimation of the bulk Ge gap is seen (yields a gap of $1.18$ eV).
 This shows the sometimes inconsistent nature of ‘one-shot’ $GW$, even with a hybrid functional starting point used in previous studies Rinke et al.
@@
 (2019).
-Finally, using linear response TDDFT  Casida (1996); Tretiak and Chernyak (2003), within the Tamm-Dancoff approximation, it is possible to obtain the linear absorption spectra of the reconstructed surfaces, here in arbitrary units for the oscillator strength.
+Finally, using linear response TDDFT  Casida (1996); Tretiak and Chernyak (2003), within the Tamm-Dancoff approximation (TDA), it is possible to obtain the linear absorption spectra of the reconstructed surfaces, here in arbitrary units for the oscillator strength.
+The TDA has been found to be accurate for systems of similar excitonic effect, compared to the full Casida equation Sander et al.
+(2015); Byun and Ullrich (2017).
 The results of our calculations are seen in Fig.
@@
 (2011) from hybrid functionals, owing to the use of TDDFT rather than ground-state DFT.
-We find that both the SRSH and HSE functionals perform with good accuracy, and get within 0.08 and and 0.14 eV of the experimental optical gap for Si respectively (for Ge the errors are 0.13 eV and 0.16 eV, respectively).
-When compared to previously reported $GW$-BSE optical gaps for the Si case, the SRSH and HSE are within 0.12 and 0.1 eV (for Ge those errors are 0.15 and 0.14 eV) respectively.
+We find that both the SRSH and HSE functionals perform with good accuracy, and get within 0.08 and and 0.13 eV of the experimental optical gap for Si respectively (for Ge the errors are 0.13 eV and 0.16 eV, respectively).
+When compared to previously reported $GW$-BSE optical gaps for the Si case, the SRSH and HSE are within 0.12 and 0.17 eV (for Ge those errors are 0.15 and 0.14 eV) respectively.
 PBE0 tends to overestimate the optical gap for the Si and Ge surface.
@@
 (2011).
-In the SISup , we report the individual excitons that make up the absorbance peaks and analyze the effects of Lorenzian broadening.
+In the SISup , we report the individual excitons that make up the absorbance peaks and analyze the effects of Lorentzian broadening.
 Other studies have shown that in-gap states induce novel low-lying excitons not present in the bulk, e.g., those originating with a mixture of bulk states and surface states Hernangómez-Pérez et al.
@@
 More importantly, we find that it is possible to obtain accurate optical gaps, using these hybrid functionals, with TDDFT calculations.
-While both SRSH and HSE functionals perform equally well here, we expect to see deviations in the performance of larger gap materials, similar to the bulk case (e.g.
+While both SRSH and HSE functionals perform equally well here, we expect to see deviations in the performance of larger gap materials, similar to the bulk case e.g.
 Fig.
 3 in SI of Wing et al.
-(2021), and Chen et al.
-(2018)).
+(2021), and Ref.
+Chen et al.
+(2018).
+Such comparisons involving HSE and PBE0 of bulk materials have also been studied e.g.
+Song et al.
+(2013), and table SIII.1 of Gant et al.
+(2022).
 Previous work more than a decade ago Jain et al.
@@
 Finally, while a tuning procedure is not yet available for surfaces, the reliable performance seen with the transferability of bulk parameters for these reconstructed surfaces, provides an indication that the construction of this class of hybrids goes in the correct direction.
-I Acknowledgements
+I Supporting Information
+Computational details, convergence tests, projected band structures for functionals not included in main text, and atomic structure details.
+II Acknowledgements
 F.S.
@@
 M.C.-G.
-is grateful to the Azrieli foundation for the award of an International Postdoctoral Fellowship.
+is grateful to the Azrieli foundation for the award of an International Postdoctoral Fellowship and acknowledges support from the Diputación Foral de Gipuzkoa through Grant 2024-FELL-000007-01.
+F.R.
+acknowledges support from the BEWARE scheme of the Wallonia-Brussels Federation for funding under the European Commission’s Marie Curie-Skłodowska Action (COFUND 847587).
 L.K.